# Parallel transport on matrix manifolds and Exponential Action

### v1

@@ 6. A quotient metric
$\begin{gathered}\dot{\Delta}=\gamma U^{-1}(\mathtt{a}W+(1+\beta)[W,\mathtt{a}_{\mathfrak{a}}]+\dot{W})U,\\ \Gamma(\dot{\gamma},\Delta)=-\frac{1}{2}\gamma U^{-1}(\mathtt{a}W+W\mathtt{a}+U[U^{-1}\mathtt{a}U,U^{-1}WU]_{\mathfrak{k}}U^{-1})U\\ +\frac{1+\beta}{2}\gamma U^{-1}\left([\mathtt{a}_{\mathfrak{a}},W]+[W_{\mathfrak{a}},\mathtt{a}]\right)U,\end{gathered}$
which simplifies to the equation for $W$. When $(U^{-1}[W,\mathtt{a}]U)_{\mathfrak{k}}=U^{-1}[W,\mathtt{a}]_{\mathfrak{k}}U$ holds, the coefficients of that equation are constant and we have eq. 41.
That simplified condition holds when $\beta=-1$, as it implies $U=\operatorname{I}_{n}$ and the rest follows. In this case, $\mathrm{P}_{\mathtt{a}}\mathtt{b}=\frac{1}{2}[\mathtt{b},\mathtt{a}]_{\mathfrak{m}}\in\mathfrak{m}$, or $\mathrm{P}_{\mathtt{a}}$ maps to $\mathfrak{m}$.
For the remaining simplified case, when $\mathfrak{k}\subset\mathfrak{a}_{{}_{\top}}$ then we have $[\mathtt{a}_{\mathfrak{a}},\mathtt{c}]_{\mathfrak{k}}=0$ for $\mathtt{c}\in\mathfrak{g}$, since set $\mathtt{c}=\mathtt{c}_{\mathfrak{a}_{{}_{\top}}}+\mathtt{c}_{\mathfrak{a}}+\mathtt{c}_{\mathfrak{a}_{{}_{\Join}}}$, the bracket reduces to $[\mathtt{a}_{\mathfrak{a}},\mathtt{c}_{\mathfrak{a}}+\mathtt{c}_{\mathfrak{a}_{{}_{\Join}}}]_{\mathfrak{k}}\subset(\mathfrak{a}+\mathfrak{a}_{{}_{\Join}})_{\mathfrak{k}}=0$. This shows $(\operatorname{ad}_{\mathtt{a}_{\mathfrak{a}}}^{j}\mathtt{c})_{\mathfrak{k}}=0$ for $j>0$ or $(U^{-1}[W,\mathtt{a}]U)_{\mathfrak{k}}=[W,\mathtt{a}]_{\mathfrak{k}}$. Similarly, $U[W,\mathtt{a}]_{\mathfrak{k}}U^{-1}=[W,\mathtt{a}]_{\mathfrak{k}}$ as $\mathtt{a}_{\mathfrak{a}}$ commutes with $\mathfrak{k}\subset\mathfrak{a}_{{}_{\top}}$, hence the simplified condition is satisfied.
To show $\mathrm{P}_{\mathtt{a}}$ operates on $\mathfrak{m}$, the previous paragraph showed $[\mathtt{a}_{\mathfrak{a}},\mathtt{b}]_{\mathfrak{k}}=0$ and $[\mathtt{b}_{\mathfrak{a}},\mathtt{a}]_{\mathfrak{k}}=0$, so all three terms of $\mathrm{P}_{\mathtt{a}}$ are in $\mathfrak{m}$.
By theorem 1, $\mathrm{P}^{\mathfrak{g}}_{\mathtt{a}}:\mathtt{b}\mapsto\frac{1}{2}\left([\mathtt{b},\mathtt{a}]+(1+\beta)([\mathtt{a}_{\mathfrak{a}},\mathtt{b}]-[\mathtt{b}_{\mathfrak{a}},\mathtt{a}])\right)$ is antisymmetric under the inner product $\langle\rangle_{\vec{\beta}}$. Thus, we only need to prove $\mathtt{b}\mapsto[\mathtt{b},\mathtt{a}]_{\mathfrak{k}}$ is antisymmetric, when restricted to $\mathfrak{m}$. But this is clear as $\mathfrak{k}$ and $\mathfrak{m}$ are orthogonal under $\langle\rangle_{\vec{\beta}}$. ∎
@@ 7.1. An isometry and a totally geodesic submanifold
In a sphere, great circles are geodesics. Given a point $Y$ on the unit sphere, and a tangent vector $\xi$, the plane spanned by $Y$ and $\xi$ cuts a great circle. Changing to the coordinate defined by $Y$ and the unit vector in direction $\xi$, this great circle maps bijectively and isometrically to the unit circle in $\mathbb{R}^{2}$. There is a similar picture for Stiefel manifolds. We start the following lemma, first discussed in [19, section 5.1].
@@
intersects $\mathrm{St}_{n,d}$ along a submanifold $\mathcal{M}_{\mathcal{S}}\subset\mathcal{M}$, which maps bijectively to $\mathrm{St}_{d+k,d}$. This bijection could be defined as
(43) $\mathfrak{F}:V\mapsto F^{\mathsf{T}}V\text{ for }F=F_{Y,Q}=\begin{bmatrix}Y|Q\end{bmatrix},\text{ with }V\in\mathcal{M}_{\mathcal{S}}.$
We have $\mathfrak{F}Y=\operatorname{I}_{d+k,d}=\begin{bmatrix}\operatorname{I}_{d}\\ O_{k\times d}\end{bmatrix}\in\mathrm{St}_{d+k,d}\subset\mathbb{R}^{n\times(d+k)}$. If $\xi=YY^{\mathsf{T}}\xi$, we can consider $F_{Y,Q}=Y$, or $Q=0$ and the intersection $\mathcal{M}_{\mathcal{S}}$ is isometric to $\operatorname{O}(d)$, with two connected components.
@@
First, if $\xi\neq YY^{\mathsf{T}}\xi$, since columns of $Q_{\xi}$ are in the column span of $\xi-YY^{\mathsf{T}}\xi$, we have $Y^{\mathsf{T}}Q_{\xi}=0$, and $Q_{\xi}^{\mathsf{T}}Q_{\xi}=\operatorname{I}_{k}$ by the orthogonal assumption. Express the columns of $\xi-YY^{\mathsf{T}}\xi$ in the basis $Q_{\xi}$ as $\xi-YY^{\mathsf{T}}\xi=Q_{\xi}R_{\xi}$ for some matrix $R_{\xi}$, then $\xi=Y(Y^{\mathsf{T}}\xi)+Q_{\xi}R_{\xi}$. This shows $Q_{\xi}$ has the desired property.
Assume $Y^{\mathsf{T}}Q=0,Q^{\mathsf{T}}Q=\operatorname{I}_{k}$ and $\xi=YA+QR$ for some matrix $A,R$. With $F=\begin{bmatrix}Y|Q\end{bmatrix}$, we have $F^{\mathsf{T}}F=\operatorname{I}_{d+k,d+k}$. If $V=YV_{Y}+QV_{Q}\in\mathcal{M}_{\mathcal{S}}$, then $\mathfrak{F}V=F^{\mathsf{T}}V=\begin{bmatrix}V_{Y}\\ V_{Q}\end{bmatrix}$. The condition $V^{\mathsf{T}}V=\operatorname{I}_{d}$ implies $V_{Y}^{\mathsf{T}}V_{Y}+V_{Q}^{\mathsf{T}}V_{Q}=\operatorname{I}_{d}$ by substitution, but this means $(\mathfrak{F}V)^{\mathsf{T}}(\mathfrak{F}V)=\operatorname{I}_{d}$. Thus, $\mathfrak{F}$ maps to $\mathrm{St}_{d+k,d}$. Furthermore, $V=F(\mathfrak{F}V)$ allows us to compute the inverse of $\mathfrak{F}$ to $\mathcal{M}_{\mathcal{S}}$, hence $\mathfrak{F}$ is a bijection.
By direct calculation, $F^{\mathsf{T}}Y=\operatorname{I}_{d+k,d}$ and $F^{\mathsf{T}}\xi=\begin{bmatrix}A\\ R\end{bmatrix}\in T_{\operatorname{I}_{d+k,d}}\mathrm{St}_{d+k,d}$.
@@
which is the metric [11], using the parametrization in [19]. For a tangent vector $\nu$ at $V\in\mathcal{M}_{\mathcal{S}}$, write $\nu=F\mathfrak{F}(\nu),V=F\mathfrak{F}(V)$ and substitute, we get
$\langle\nu,\nu\rangle_{\vec{\beta}}=\operatorname{Tr}\mathfrak{F}(\nu)^{\mathsf{T}}\mathfrak{F}(\nu)+(\alpha-1)\operatorname{Tr}\mathfrak{F}(\nu)^{\mathsf{T}}\mathfrak{F}(V)\mathfrak{F}(V)^{\mathsf{T}}\mathfrak{F}(\nu).$
Thus, if we equip $\mathrm{St}_{d+k,d}$ with a metric with the same parameter $\alpha$, $\mathfrak{F}$ is an isometry.
Computation done on this smaller Stiefel manifold is more effective, as observed in the case of geodesics [5, 19, 31]. The case of parallel transport will be done similarly.
By [23, page 104], $\mathcal{M}_{\mathcal{S}}$ is a totally geodesic submanifold, as the geodesic with initial point $V\in\mathcal{M}_{\mathcal{S}}$ and initial vector tangent to $\mathcal{M}_{\mathcal{S}}$ will stay on $\mathcal{M}_{\mathcal{S}}$.
@@ 7.2. Submersion metric, connection and parallel transport
The Stiefel manifold $\mathrm{St}_{n,d}$ could be consider as a quotient manifold of $\mathrm{SO}(n)$ under the map sending an $n\times n$ matrix to its first $d$ columns, or $\mathfrak{q}:X\mapsto X\operatorname{I}_{n,d}$, for $X\in\mathrm{SO}(n)$, $\operatorname{I}_{n,d}=\begin{bmatrix}\operatorname{I}_{d}\\ 0_{(n-d)\times d}\end{bmatrix}$. This map is onto, realizing $\mathrm{St}_{n,d}$ as a quotient of $\mathrm{SO}(n)$ by the group $\mathrm{K}$ of block diagonal matrices of the form $\operatorname{diag}(\operatorname{I}_{d},Q)$ where $Q\in\mathrm{SO}(n-d)$. This is due to the fact that we can complement an orthogonal matrix $Y$ with an orthogonal matrix $Y_{\perp}\in\mathrm{St}_{n,n-d}$ to an element of $\mathrm{SO}(n)$, and given $Y_{\perp}$, any other complement basis is of the form $Y_{\perp}Q$ for $Q\in\mathrm{K}$. In this case, $\mathfrak{a}$ is the top $d\times d$ block-matrix copy of $\mathfrak{so}(d)$ and $\mathfrak{k}=\mathfrak{a}_{{}_{\top}}$ is the bottom $(n-d)\times(n-d)$ copy of $\mathfrak{so}(n-d)$. The complement $\mathfrak{m}$ consists of all matrices with the lower right $(n-d)\times(n-d)$ blocks identically zero. Thus, graphically, $\mathfrak{k}=\begin{bmatrix}0_{d\times d}&0\\ 0&\mathfrak{so}(n-d)\end{bmatrix},\mathfrak{m}=\begin{bmatrix}\mathfrak{a}&\mathfrak{a}_{{}_{\Join}}\\ \mathfrak{a}_{{}_{\Join}}&0_{(n-d)\times(n-d)}\end{bmatrix}$.
The differential $d\mathfrak{q}$ at $\operatorname{I}_{n}$ also maps a Lie algebra element in $\mathfrak{so}(n)$ to its first $d$ columns. If $\Delta=\begin{bmatrix}A\\ B\end{bmatrix}$ is a tangent vector at $\operatorname{I}_{n,d}$, then its horizontal lift is the matrix $\bar{\Delta}=\begin{bmatrix}A&-B^{\mathsf{T}}\\ B&0\end{bmatrix}$, the bottom right entry corresponding to $\mathfrak{k}$ is set to zero. With $\beta_{0}=-\frac{1}{2},\beta_{1}=\alpha,\beta=\frac{\beta_{1}}{\beta_{0}}=-2\alpha$, the metric in eq. 33 at $\operatorname{I}_{n}$ is
@@
which is antisymmetric, and $\operatorname{I}_{{}_{\perp}}^{\mathsf{T}}(X^{\mathsf{T}}\bar{\xi})\operatorname{I}_{{}_{\perp}}=0$. We also have $(X^{\mathsf{T}}\bar{\xi})_{\mathfrak{a}}=\operatorname{I}_{n,d}Y^{\mathsf{T}}\xi\operatorname{I}_{n,d}$.
Under this metric and with the lift above, using eq. 38, we can show [11, 20] the Levi-Civita connection for Stiefel manifold is $\nabla_{\mathtt{Z}}\mathtt{W}=\mathrm{D}_{\mathtt{Z}}\mathtt{W}+\Gamma(\mathtt{Z},\mathtt{W})$ with the Christofel function evaluated at $Y\in\mathcal{M}=\mathrm{St}_{n,d},\mathtt{Z}(Y)=\xi,\mathtt{W}(Y)=\eta$ is

### v2

@@ 7. A naturally reductive family of quotient metrics
(49) $\begin{gathered}\Gamma^{\mathcal{H}}(\xi,\eta)=-\frac{1}{2}\left(\xi X^{-1}\eta+\eta X^{-1}\xi+X(\mathcal{P}_{\mathfrak{k}}[X^{-1}\xi,X^{-1}\eta])\right)\\ +\frac{1}{2}X\left([(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}(X^{-1}\xi),X^{-1}\eta]+[(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}(X^{-1}\eta),X^{-1}\xi]\right)\end{gathered}$
if we express the horizontal connection for two horizontal vector fields $\mathtt{Y},\mathtt{Z}$ as $\nabla^{\mathcal{H}}_{\mathtt{Y}}\mathtt{Z}=\mathrm{D}_{\mathtt{X}}\mathtt{Y}+\Gamma^{\mathcal{H}}(\mathtt{X},\mathtt{Y})$. The horizontal geodesic $\gamma$ with $\gamma(0)=X,\dot{\gamma}(0)=\xi$, where $\xi$ is a horizontal vector is given by eq. 28.
Let $\mathtt{V}=X^{-1}\xi\in\mathfrak{m}$. The parallel transport of $\eta\in\mathcal{H}_{X}$ is given by
(50) $\mathtt{T}(\gamma)_{0}^{t}\eta=X\operatorname{expm}(t\left(\mathtt{V}-(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\mathtt{V}\right))\operatorname{expv}(t\mathrm{P}_{\mathtt{V}},X^{-1}\eta)\operatorname{expm}(t(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\mathtt{V})$
(51)  $\text{ where }\mathrm{P}_{\mathtt{V}}:\mathtt{B}\mapsto\frac{1}{2}\left(\mathcal{P}_{\mathfrak{m}}[\mathtt{B},\mathtt{V}]+([(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\mathtt{V},\mathtt{B}]-[(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\mathtt{B},\mathtt{V}])\right).$
$\mathrm{P}_{\mathtt{V}}$ maps $\mathfrak{m}$ to itself and the operator $\mathrm{P}_{\mathtt{V}}$ is antisymmetric in the pairing $\langle\rangle_{\vec{\beta}}$ on $\mathfrak{m}$.
Proof.
The proof is similar to the proof of theorem 2. We will use the notations and results from that proof.
We will assume $X=\operatorname{I}_{\mathcal{G}}$ then use left invariance. For two horizontal vectors $\xi_{0},\eta_{0}\in\mathfrak{m}$ at $X=\operatorname{I}_{\mathcal{G}}$, consider the horizontal vector fields $\bar{\xi}:X\mapsto X\xi_{0},\bar{\eta}:X\mapsto X\eta_{0}$, then by [33, Lemma 7.45], if $\mathcal{H}$ is the horizontal projection and $\Gamma$ is the Christoffel function of Levi-Civita connection in eq. 27 then
$\begin{gathered}\Gamma^{\mathcal{H}}_{X=I_{\mathcal{G}}}(\bar{\xi},\bar{\eta})=\mathcal{H}(\mathrm{D}_{\bar{\xi}}\bar{\eta}+\Gamma(\bar{\xi},\bar{\eta}))-\mathrm{D}_{\bar{\xi}}\bar{\eta}\\ =\mathcal{P}_{\mathfrak{m}}(\xi_{0}\eta_{0}-\frac{1}{2}(\xi_{0}\eta_{0}+\eta_{0}\xi_{0}))+\frac{1}{2}\left([(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\xi_{0},\eta_{0}]+[(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\eta_{0},\xi_{0}]\right)-\xi_{0}\eta_{0}\end{gathered}$
because by lemma 3, the sum of the two Lie brackets in the third and fourth terms above are in $\mathfrak{m}$. We collect the remaining terms to
$\begin{gathered}\frac{1}{2}\mathcal{P}_{\mathfrak{m}}[\xi_{0},\eta_{0}]-\xi_{0}\eta_{0}=\frac{1}{2}[\xi_{0},\eta_{0}]-\frac{1}{2}\mathcal{P}_{\mathfrak{k}}[\xi_{0},\eta_{0}]-\xi_{0}\eta_{0}=-\frac{1}{2}(\xi_{0}\eta_{0}+\eta_{0}\xi_{0})-\frac{1}{2}\mathcal{P}_{\mathfrak{k}}[\xi_{0},\eta_{0}].\end{gathered}$
Together, we get $\Gamma^{\mathcal{H}}_{X=I_{\mathcal{G}}}(\bar{\xi},\bar{\eta})$ evaluated at $X=\operatorname{I}_{\mathcal{G}}$ is given by eq. 49. We get the general case by using left invariance.
The result on the horizontal lift of geodesics is standard. For parallel transport, we will assume again $X=\operatorname{I}_{\mathcal{G}}$, the general result will follow by left invariance. As before, by theorem 3 and [39, Proposition 2.12], the lifted parallel transport from $(\operatorname{I}_{\mathcal{G}},\operatorname{I}_{\mathcal{A}})$ is $(\exp(t(\xi+\xi^{\mathcal{A}}))\mathtt{C}^{\mathcal{G}},\exp(t\xi^{\mathcal{A}})\mathtt{C}^{\mathcal{A}})$ where $(\mathtt{C}^{\mathcal{G}}(t),\mathtt{C}^{\mathcal{A}}(t))\in\tilde{\mathfrak{m}}_{\bar{\mathfrak{q}}}$ solves
$\begin{gathered}(\dot{\mathtt{C}}^{\mathcal{G}},\dot{\mathtt{C}}^{\mathcal{A}})=\mathcal{P}_{\tilde{\mathfrak{m}}_{\bar{\mathfrak{q}}}}(-\frac{1}{2}[\xi+\xi^{\mathcal{A}},\mathtt{C}^{\mathcal{G}}],-\frac{1}{2}[\xi^{\mathcal{A}},\mathtt{C}^{\mathcal{A}}]),\\ (\mathtt{C}^{\mathcal{G}}(0),\mathtt{C}^{\mathcal{A}}(0))=(\eta+\eta^{\mathcal{A}},\eta^{\mathcal{A}}).\end{gathered}$
The lifted transport maps to $\exp(t(\xi+\xi^{\mathcal{A}}))(\mathtt{C}^{\mathcal{G}}-\mathtt{C}^{\mathcal{A}})\exp(-t\xi^{\mathcal{A}})$, the horizontal transport on $\mathcal{G}$, and $F(t)=\mathtt{C}^{\mathcal{G}}(t)-\mathtt{C}^{\mathcal{A}}(t)$ solves
$\begin{gathered}\dot{F}=-\frac{1}{2}\mathcal{P}_{\mathfrak{m}}\left([\xi,F]-[\xi,(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}F]-[(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\xi,F]\right).\end{gathered}$
The additional projection $\mathcal{P}_{\mathfrak{m}}$ is from the map $d\pi$ in $d\bar{\mathfrak{q}}=d\pi\circ d\mathfrak{q}$. Thus, $F(t)=\operatorname{expv}(t\mathrm{P}_{\xi},\eta)$, giving us the parallel transport formula eq. 50 for $X=\operatorname{I}_{\mathcal{G}}$, where we note that the sum of the last two brackets are in $\mathfrak{m}$ by lemma 3. The remaining statements are immediate or follow from similar arguments in the proof of theorem 2. ∎
@@ 8.1. An isometry and a totally geodesic submanifold
In a sphere, great circles are geodesics. Given a point $Y$ on the unit sphere, and a tangent vector $\xi$, the plane spanned by $Y$ and $\xi$ cuts a great circle. Changing to the coordinate defined by $Y$ and the unit vector in the direction $\xi$, this great circle maps isometrically to the unit circle in $\mathbb{R}^{2}$. There is a similar picture for Stiefel manifolds. We start the following lemma, with ideas in [25] and discussed partially in [29, section 5.1].
@@
intersects $\mathrm{St}_{n,d}$ along a submanifold $\mathcal{M}_{\mathcal{S}}\subset\mathcal{M}$. Let $\mathsf{F}$ be the map from $\mathbb{R}^{n\times d}$ to $\mathbb{R}^{(d+k)\times d}$ defined by
(53) $\mathsf{F}:V\mapsto FV,\quad V\in\mathbb{R}^{n\times d}\quad\text{ for }F=F_{Y,Q}=\begin{bmatrix}Y^{\mathsf{T}}\\ Q^{\mathsf{T}}\end{bmatrix}$.
Then $\mathsf{F}$ restricted to $\mathcal{M}_{\mathcal{S}}$ is a bijection onto $\mathrm{St}_{d+k,d}\subset\mathbb{R}^{(d+k)\times d}$. We have $\mathsf{F}Y=\operatorname{I}_{d+k,d}=\begin{bmatrix}\operatorname{I}_{d}\\ O_{k\times d}\end{bmatrix}\in\mathrm{St}_{d+k,d}\subset\mathbb{R}^{n\times(d+k)}$. If $\xi=YY^{\mathsf{T}}\xi$, we can consider $F_{Y,Q}=Y^{\mathsf{T}}$, or $Q=0$ and the intersection $\mathcal{M}_{\mathcal{S}}$ is isometric to $\operatorname{O}(d)$, with two connected components.
@@
First, if $\xi\neq YY^{\mathsf{T}}\xi$, since columns of $Q_{\xi}$ are in the column span of $\xi-YY^{\mathsf{T}}\xi$, we have $Y^{\mathsf{T}}Q_{\xi}=0$, and $Q_{\xi}^{\mathsf{T}}Q_{\xi}=\operatorname{I}_{k}$ by the orthonormal assumption. Express the columns of $\xi-YY^{\mathsf{T}}\xi$ in the basis $Q_{\xi}$ as $\xi-YY^{\mathsf{T}}\xi=Q_{\xi}R_{\xi}$ for some matrix $R_{\xi}$, then $\xi=Y(Y^{\mathsf{T}}\xi)+Q_{\xi}R_{\xi}$. This shows $Q_{\xi}$ has the desired property.
Assume $Y^{\mathsf{T}}Q=0,Q^{\mathsf{T}}Q=\operatorname{I}_{k}$ and $\xi=YA+QR$ for some matrix $A,R$. With $F=\begin{bmatrix}Y|Q\end{bmatrix}^{\mathsf{T}}$, we have $FF^{\mathsf{T}}=\operatorname{I}_{d+k,d+k}$. If $V=YV_{Y}+QV_{Q}\in\mathcal{M}_{\mathcal{S}}$, then $\mathsf{F}V=FV=\begin{bmatrix}V_{Y}\\ V_{Q}\end{bmatrix}$. The condition $V^{\mathsf{T}}V=\operatorname{I}_{d}$ implies $V_{Y}^{\mathsf{T}}V_{Y}+V_{Q}^{\mathsf{T}}V_{Q}=\operatorname{I}_{d}$ by substitution, but this means $(\mathsf{F}V)^{\mathsf{T}}(\mathsf{F}V)=\operatorname{I}_{d}$. Thus, $\mathsf{F}$ maps $\mathcal{M}_{\mathcal{S}}$ to $\mathrm{St}_{d+k,d}$. Conversely, $V=F^{\mathsf{T}}(\mathsf{F}V)$ and the map $W\mapsto F^{\mathsf{T}}W$ maps $\mathrm{St}_{d+k,d}$ to $\mathcal{M}_{\mathcal{S}}$, hence, $\mathsf{F}$ is a bijection from $\mathcal{M}_{\mathcal{S}}$ to $\mathrm{St}_{d+k,d}$.
By direct calculation, $FY=\operatorname{I}_{d+k,d}$ and $F\xi=\begin{bmatrix}A\\ R\end{bmatrix}\in T_{\operatorname{I}_{d+k,d}}\mathrm{St}_{d+k,d}$.
@@
which is the metric [16], using the parametrization in [29]. For a tangent vector $\nu$ at $V\in\mathcal{M}_{\mathcal{S}}$, write $\nu=F^{\mathsf{T}}\mathsf{F}(\nu),V=F^{\mathsf{T}}\mathsf{F}(V)$ and substitute, we get
$\langle\nu,\nu\rangle_{\vec{\beta}}=\operatorname{Tr}\mathsf{F}(\nu)^{\mathsf{T}}\mathsf{F}(\nu)+(\alpha-1)\operatorname{Tr}\mathsf{F}(\nu)^{\mathsf{T}}\mathsf{F}(V)\mathsf{F}(V)^{\mathsf{T}}\mathsf{F}(\nu).$
Thus, if we equip $\mathrm{St}_{d+k,d}$ with a metric with the same parameter $\alpha$, $\mathsf{F}$ is an isometry.
Computing on this smaller Stiefel manifold is more effective, as observed in the case of geodesics [7, 29, 46, 35, 45]. The case of parallel transport will be done similarly. Besides the original paper [7], the first author benefited from the prior works [35, Chapter 5] and [45] in understanding the advantage of working on the baby problem, as referred to in the first citation.
By [33, page 104], $\mathcal{M}_{\mathcal{S}}$ is a totally geodesic submanifold, as the geodesic with initial point $V\in\mathcal{M}_{\mathcal{S}}$ and initial vector tangent to $\mathcal{M}_{\mathcal{S}}$ will stay on $\mathcal{M}_{\mathcal{S}}$.
@@ 8.2. Submersion metric, connection and parallel transport
The material on the submersion metric and the connection of this section follows from the treatment of Stiefel manifolds in the original paper [7], the thesis [35] (both with $\alpha\in\{\frac{1}{2},1\}$), and [19, 16], where the quotient structure $(\mathcal{G}\times\mathcal{A})/\tilde{\mathcal{K}}$ with $\mathcal{G}=\mathrm{SO}(n),\mathcal{A}=\mathrm{SO}(d),\mathcal{K}=\mathrm{SO}(n-d)$ and $\tilde{\mathcal{K}}$ as in theorem 3 is covered with a different metric parametrization.
The Stiefel manifold $\mathrm{St}_{n,d}$ could be considered as a quotient manifold of $\mathrm{SO}(n)$ under the map sending an $n\times n$ matrix to its first $d$ columns, or $\Theta:X\mapsto X\operatorname{I}_{n,d}$, for $X\in\mathrm{SO}(n)$, $\operatorname{I}_{n,d}=\begin{bmatrix}\operatorname{I}_{d}\\ 0_{(n-d)\times d}\end{bmatrix}$. This map is surjective, realizing $\mathrm{St}_{n,d}$ as a quotient of $\mathrm{SO}(n)$ by the group $\mathcal{K}$ of block diagonal matrices of the form $\operatorname{diag}(\operatorname{I}_{d},Q)$ where $Q\in\mathrm{SO}(n-d)$. This is due to the fact that we can complement an orthogonal matrix $Y$ with an orthogonal matrix $Y_{\perp}\in\mathrm{St}_{n,n-d}$ to an element of $\mathrm{SO}(n)$, and given $Y_{\perp}$, any other complement basis is of the form $Y_{\perp}Q$ for $Q\in\mathcal{K}$. In this case, $\mathfrak{a}$ is the top $d\times d$ block-matrix copy of $\mathfrak{so}(d)$ and $\mathfrak{k}$ is the bottom $(n-d)\times(n-d)$ copy of $\mathfrak{so}(n-d)$. The complement $\mathfrak{m}$ consists of all matrices with the lower right $(n-d)\times(n-d)$ blocks identically zero. Thus, graphically, $\mathfrak{k}=\begin{bmatrix}0_{d\times d}&0\\ 0&\mathfrak{so}(n-d)\end{bmatrix},\mathfrak{m}=\begin{bmatrix}\mathfrak{a}&*\\ *&0_{(n-d)\times(n-d)}\end{bmatrix}$.
The differential $d\Theta$ at $\operatorname{I}_{n}$ also maps a Lie algebra element in $\mathfrak{so}(n)$ to its first $d$ columns. If $\Delta=\begin{bmatrix}A\\ B\end{bmatrix}$ is a tangent vector at $\operatorname{I}_{n,d}$, then its horizontal lift is the matrix $\bar{\Delta}=\begin{bmatrix}A&-B^{\mathsf{T}}\\ B&0\end{bmatrix}$, the bottom right entry corresponding to $\mathfrak{k}$ is set to zero. With $\beta_{0}=-\frac{1}{2},\mathcal{B}=\alpha\operatorname{I}_{\mathfrak{a}}$, the metric in eq. 39 at $\operatorname{I}_{n}$ is
@@
which is antisymmetric, and $\operatorname{I}_{{}_{\perp}}^{\mathsf{T}}(X^{\mathsf{T}}\bar{\xi})\operatorname{I}_{{}_{\perp}}=0$. We also have $\mathcal{P}_{\mathfrak{a}}(X^{\mathsf{T}}\bar{\xi})=\operatorname{I}_{n,d}Y^{\mathsf{T}}\xi\operatorname{I}_{n,d}^{\mathsf{T}}$.
Under this metric and with the lift above, using eq. 49, we can show [16, 30] the Levi-Civita connection for Stiefel manifold is $\nabla_{\mathtt{Z}}\mathtt{W}=\mathrm{D}_{\mathtt{Z}}\mathtt{W}+\Gamma(\mathtt{Z},\mathtt{W})$ with the Christoffel function evaluated at $Y\in\mathcal{M}=\mathrm{St}_{n,d},\mathtt{Z}(Y)=\xi,\mathtt{W}(Y)=\eta$ is